# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

### v2

@@ 6 Example: Behrens–Fisher problem
For brevity, I only give one detailed illustration of the inner probabilistic approximation here, but two additional examples are presented in Appendix F.
Few examples have deeper practical and historical relevance than the Behrens–Fisher problem, whose name comes from Fisher’s reference to a technical result of Walter Behrens in his proposed fiducial solution to the problem (Fisher, 1935a ; Fisher, 1939). The original problem is simple to state: independent samples of size $n_{1}$ and $n_{2}$, respectively, are available from two distinct normal populations, ${\sf N}(\Theta_{11},\Theta_{12}^{2})$ and ${\sf N}(\Theta_{21},\Theta_{22}^{2})$, with $\Theta=(\Theta_{11},\Theta_{12},\Theta_{21},\Theta_{22})$ unknown, and the goal is marginal inference on the difference $\Phi=m(\Theta)=\Theta_{21}-\Theta_{11}$ of the two means. The problem is straightforward if the two variances are known or if their ratio is known, but the fully unknown variance case has remained elusive, i.e., lots of candidate solutions are available but, despite its apparent simplicity and practical relevance, there is no consensus on which solution is “right” or “best.” For a review of the problem and various solutions, see Kim and Cohen, (1998).
By far the most widely used solution to the Behrens–Fisher problem is the modification of the basic Student-t pivot with the degrees of freedom approximations due to Welch, (1938, 1947); this is implemented in R’s t.test function. Other standard approaches include the simple-but-conservative solution proposed by Hsu, (1938) and Scheffé, (1970), and the Bayesian solution proposed by Jeffreys, (1940), based on the right Haar prior for $\Theta$, which is mathematically equivalent to Fisher’s fiducial solution. Ironically, the solution proposed by Jeffreys is different from the Bayesian solution based on the Jeffreys prior, both in its construction and its performance. A conservative IM solution was advanced in Martin and Liu, 2015b , but here I take a different approach.
@@
The new approach I’m proposing here first carries out a different—and generally more efficient—IM marginalization strategy, based on validifying the profile relative likelihood. This idea was first developed in Martin, 2023b and an illustration of its use in the Behrens–Fisher problem is considered in Example 5 there. Roughly speaking, the difference between this profile-based marginal IM construction and the aforementioned extension-based construction is where the marginalization is carried out: the former first eliminates nuisance parameters in the relative likelihood and then constructs the IM contour directly for $\Phi$, whereas the latter first constructs the IM contour for $\Theta$ and then marginalizes to $\Phi$. It had been conjectured that the profile-based marginal IM construction was more efficient than the extension-based construction, and Martin and Williams, (2025) recently confirmed this, at least asymptotically. My proposal here is to extract the inner probabilistic approximation of this profile-based marginal possibilistic IM for $\Phi$. This requires no change in the computational approach outlined in Section 5, and the exact validity of the marginal IM implies, e.g., that credible intervals for $\Phi$ based on this “posterior” are exact confidence intervals. The only wrinkle is that the relative profile likelihood doesn’t have a closed-form expression in the Behrens–Fisher problem, and its distribution depends on a nuisance parameter; this is precisely what makes the Behrens–Fisher problem challenging. This increases the computational cost of the inner probabilistic approximation compared to other examples (see Appendix F), but this pays off in terms of efficiency gains as I demonstrate below.
@@
Method Coverage probability Mean length
@@ 7 Conclusion
This paper offers a new perspective on no-prior Bayesian inference, one that has certain close ties to Bayesian inference but is not actually Bayesian at all. My approach starts with an inferential model (IM) framework for data-driven uncertainty quantification that prioritizes reliability, insisting on calibration of its data-dependent degrees of belief. The kind of calibration IMs require makes them incompatible with probabilistic Bayesian inference, and it’s for this reason that the IM’s output possibilistic, i.e., takes the mathematical form of a possibility measure. However, if one desires probabilistic uncertainty quantification, then the IM can accommodate this with an inner probabilistic approximation of its possibilistic output. Aside from being exactly probability matching, the proposed solution agrees with existing no-prior Bayes/fiducial solutions in group invariant problems where there’s an overall consensus, and is also asymptotically efficient by a version of the celebrated Bernstein–von Mises theorem. And there’s literally no choice of prior distribution required: for the given model and data, the possibilistic IM is what it is, so it’s just a matter of extracting the inner probabilistic approximation.
The characterization of the contents of the IM’s credal set in Theorem 1 suggests a strategy for evaluating, at least approximately/conservatively, the inner probabilistic approximation via Monte Carlo. I explored this thoroughly in Martin, (2025), but mostly in the context of possibilistic IMs. Here I use this same strategy for my proposed reIMagined no-prior Bayesian inference, and I show that this leads to a new and broadly reliable solution to the technically challenging and practically relevant Behrens–Fisher problem. Nothing about this approach was tailored to the Behrens–Fisher problem specifically, so I fully expect similarly good performance in many other important applications. Surely, the computational efficiency of my proposed solution can be improved, so I welcome advancements along these lines from computationally savvy readers.
The focus here has been on no-prior solutions, but there are cases in which incomplete or partial prior information is available, including high-dimensional problems in which structural assumptions like sparsity are common. Bayesian solutions can’t work with incomplete prior information—a prior distribution, perhaps diffuse, must be specified for every aspect of the unknown $\Theta$. Genuine partial-prior possibilistic IMs have recently been developed (e.g., Martin, 2022b ; Martin, 2023b ), and an interesting idea would be to apply the same inner probabilistic approximation proposed here in the no-prior case to the partial-prior case in the above references. Just as in this paper, these inner probabilistic approximations would enjoy some of the inherent reliability properties satisfied by the partial-prior IM, but the details remain to be investigated.
@@ Acknowledgments
This work is partially supported by the U.S. National Science Foundation, under grants SES–2051225 and DMS–2412628.
@@
$\sup_{\Theta\in\mathbb{T}}\mathsf{P}_{\Theta}\{\overline{\Pi}_{X}(H)\leq\alpha\text{ for some $H$ with $H\ni\Theta$}\}\leq\alpha,\quad\alpha\in[0,1].$ (17)
@@
There are mathematical obstacles that would make it difficult for any non-possibilistic quantification of uncertainty to satisfy the strong validity property in Equation (4) of the main paper, and hence the uniform calibration property in (17) above. Indeed, if $(\underline{\Omega}_{x},\overline{\Omega}_{x})$ was some other imprecise-probabilistic quantification of uncertainty about $\Theta$, given $X=x$, then one could define the analogue of its “contour,” i.e., $\theta\mapsto\overline{\Omega}_{x}(\{\theta\})$, but this may not meet the requirement of a possibility contour, that is, generally,
$\sup_{\theta\in\mathbb{T}}\overline{\Omega}_{x}(\{\theta\})<1,\quad\text{and often $\ll 1$}.$
Nothing in imprecise probability theory says that this “contour” must satisfy this supremum-equals-1 property and, in fact, in many (non-statistical) cases there are good reasons why such a property would be undesirable. But, as Shafer, (1976, p. 226) explains, statistical inference problems are special and well-suited for this supremum-equals-1 property. What I can add to Shafer’s justification is that it is virtually impossible for anything other than a possibilistic quantification of uncertainty, one for which the supremum-equals-1 property holds, to satisfy the validity property as stated in Equation (4) of the main paper and, therefore, the equivalent uniform calibration property in (17) above. The point is that, if $\overline{\Omega}_{X}(\{\Theta\})$ is bounded away from 1 for almost all $X$, then it can’t be stochastically no less than ${\sf Unif}(0,1)$. Moreover, if one was able to find a non-possibilistic IM $(\underline{\Omega}_{x},\overline{\Omega}_{x})$ that does satisfy the validity property in Equation (4) in the main text, then there is a possibilistic IM that is valid and no less efficient. So, there is at least no loss of generality in focusing on IMs that take on the possibilistic form.
@@
B.1 Theory
@@
Recall that $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ is a family of probability distributions supported on $\mathbb{X}$ that is also invariant with respect to a group of transformations $\mathscr{G}$ from $\mathbb{X}$ to itself. If the distributions $\mathsf{P}_{\theta}$ have densities with respect to some underlying dominating measure, which I’ll assume here, then invariance means
@@
where $\chi(g)$ is the “multiplier,” a change-of-variables Jacobian term; see Schervish, (1995, Ch. 6) and Eaton, (1989) for details. A further simplification that I’ll adopt here, common in the literature (e.g., Schervish, 1995, p. 371), is to assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$.
@@
Let $\{p_{\theta}:\theta\in\mathbb{T}\}$ be a family of densities invariant with respect to a locally compact topological group $\mathscr{G}$ in the sense of (18) and, for simplicity, assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$. In addition, assume that the following hold:
@@
A couple quick remarks are in order. First, for A1, existence and uniqueness of the left and associated right Haar measures is a classical result; see Halmos, (1950) and Nachbin, (1965). For A3, existence of a joint density with respect to a product measure ensures that there’s no difficulty in defining a conditional distribution for $G$, given $U=u$, which is needed below. Finally, $U=U_{X}$, as a function of $X$, is an ancillary statistic.
The simplest example is a location model where $\mathscr{G}=\overline{\mathscr{G}}=(\mathbb{R},+)$. This is an abelian group, so the left and right Haar measures are the same and both equal to Lebesgue measure. The function $x\mapsto t(x)$ in A2 consists of two components: in its “$g_{x}$” coordinate an equivariant function of $x$ that estimates the location and, in its “$u_{x}$” component, an invariant function of $x$, such as residuals. For example, $g_{x}=\bar{x}$ the arithmetic mean of $x=(x_{1},\ldots,x_{n})$ and $u_{x}=\{x_{i}-\bar{x}:i=1,\ldots,n\}$. Note that the $u_{x}$ coordinate satisfies a constraint, so, after it’s represented in a suitable lower-dimensional space $\mathbb{U}$, $\mu$ can be taken as Lebesgue measure there. There are many other problems that fit this general form; see, e.g., Chapters 1–2 of Fraser, (1968), including the exercises.
@@
  • Lemma 6.65 in Schervish, (1995) gives a Bayesian posterior distribution $\mathsf{Q}_{x}^{\text{\sc b}}$ for $\Theta$ that has the following density with respect to the right Haar prior measure:
    $q_{x}^{\text{\sc b}}(\theta)\propto f(\theta^{-1}\circ g,u),\quad\theta\in\mathbb{T},\quad x\equiv(g,u).$ (20)
@@
These points are crucial to the following calculation. Indeed, since the $U$ component of $X$ is ancillary, the minimum complexity principle in Martin, 2022b says the IM construction should condition on the observation $U=u$. Write $\pi_{g|u}(\theta)$ for the possibilistic IM’s contour, where the subscript is meant to emphasize that the conditional distribution of $X\equiv(G,U)$, given the observed value of $u$, is used in the validification step (2). Then the IM contour is defined as
$\pi_{g|u}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\mid U=u\},\quad\theta\in\mathbb{T}.$
Then the following sequence of identities are consequences of the two results above that link relative likelihood and the posterior density:
@@
$=\mathsf{Q}_{x}^{\text{\sc b}}\{f(H,u)\leq f(\theta^{-1}\circ g,u)\}$
$=\mathsf{Q}_{x}^{\text{\sc b}}\{q_{x}^{\text{\sc b}}(\Theta)\leq q_{x}^{\text{\sc b}}(\theta)\},$
@@
where the first line is by simplification of the definition of $\pi_{g|u}$; the second line is by the fact that $H:=\theta^{-1}\circ G$ is a pivot with respect to the conditional distribution of $G$, given $U=u$, under $\mathsf{P}_{\theta}$ (e.g., Schervish, 1995, Cor. 6.67); the third line is by the fact that $H:=\Theta^{-1}\circ g$ is also a pivot with respect to the Bayes posterior $\Theta\sim\mathsf{Q}_{x}^{\text{\sc b}}$ and has the same distribution as $H$ in previous line; and the last line is by simplification via (20).

### v3

@@ 6 Example: Behrens–Fisher problem
For brevity, I only give one detailed illustration of the inner probabilistic approximation here, but two additional examples are presented in Appendix G.
Few examples have deeper practical and historical relevance than the Behrens–Fisher problem (Fisher, 1935a ; Fisher, 1939). The original problem is simple to state: independent samples of size $n_{1}$ and $n_{2}$, respectively, are available from two distinct normal populations, ${\sf N}(\Theta_{11},\Theta_{12}^{2})$ and ${\sf N}(\Theta_{21},\Theta_{22}^{2})$, with $\Theta=(\Theta_{11},\Theta_{12},\Theta_{21},\Theta_{22})$ unknown, and the goal is marginal inference on the difference $\Phi=m(\Theta)=\Theta_{21}-\Theta_{11}$ of the two means. The problem is straightforward if the two variances are known or if their ratio is known, but the fully unknown variance case has remained elusive, i.e., lots of candidate solutions are available but, despite its apparent simplicity and practical relevance, there is no consensus on which solution is “right” or “best.” For a review of the problem and various solutions, see Kim and Cohen, (1998).
By far the most widely used solution to the Behrens–Fisher problem is the modification of the basic Student-t pivot with the degrees of freedom approximations due to Welch, (1938, 1947); this is implemented in R’s t.test function. Other standard approaches include the simple-but-conservative solution proposed by Hsu, (1938) and Scheffé, (1970), and the Bayesian solution proposed by Jeffreys, (1940), based on the right Haar prior for $\Theta$, which is mathematically equivalent to Fisher’s fiducial solution. Ironically, the solution proposed by Jeffreys is different from the Bayesian solution based on the Jeffreys prior, both in its construction and its performance.
@@
The new approach I’m proposing here first carries out a different—and generally more efficient—IM marginalization strategy, based on validifying the profile relative likelihood. This idea was first developed in Martin, 2023b and an illustration of its use in the Behrens–Fisher problem is considered in Example 5 there. Roughly speaking, the difference between this profile likelihood-based marginal IM construction and the aforementioned extension-based construction is where the marginalization is carried out: the former first eliminates nuisance parameters in the relative likelihood and then constructs the IM contour directly for $\Phi$, whereas the latter first constructs the IM contour for $\Theta$ and then marginalizes to $\Phi$. Martin and Williams, (2025) recently confirmed that, at least asymptotically, the profile likelihood-based marginal IM construction was more efficient than the extension-based construction. My proposal here is to extract the inner probabilistic approximation of this profile likelihood-based marginal possibilistic IM for $\Phi$. This requires no change in the computational approach outlined in Section 5, and the exact validity of the marginal IM implies, e.g., that credible intervals for $\Phi$ based on this “posterior” are exact confidence intervals. The only wrinkle is that the relative profile likelihood doesn’t have a closed-form expression, and its distribution depends on a nuisance parameter; this is precisely what makes the Behrens–Fisher problem challenging. This increases the computational cost of the inner probabilistic approximation compared to other examples (see Appendix G), but this pays off in terms of efficiency gains as I demonstrate below.
@@
Method Coverage Length
@@ 7 Conclusion
This paper offers a new perspective on no-prior Bayes inference, one that has certain close ties to Bayesian inference but is not actually Bayesian at all. My approach starts with an inferential model (IM) framework for data-driven uncertainty quantification that prioritizes reliability, insisting on calibration of its data-dependent degrees of belief. The kind of calibration IMs require makes them incompatible with probabilistic Bayesian inference, and it’s for this reason that the IM’s output is possibilistic, i.e., takes the mathematical form of a possibility measure. However, if one desires probabilistic uncertainty quantification, then the IM can accommodate this with an inner probabilistic approximation of its possibilistic output. Aside from being exactly probability matching, the proposed solution agrees with existing no-prior Bayes solutions in group equivariant problems where there’s an overall consensus, and is also asymptotically efficient by a version of the celebrated Bernstein–von Mises theorem. And there’s literally no choice of prior distribution required: for the given model and data, the possibilistic IM is what it is, so it’s just a matter of extracting a suitable inner probabilistic approximation.
The characterization of the contents of the IM’s credal set in Theorem 1 suggests a strategy for evaluating, at least approximately, the inner probabilistic approximation via Monte Carlo. I explored this thoroughly in Martin, 2025a , but mostly in the context of possibilistic IMs. Here I use this same strategy for my proposed reIMagined no-prior Bayesian inference, and I show that this leads to a new and broadly reliable solution to the technically challenging and practically relevant Behrens–Fisher problem. Nothing about this approach was tailored to the Behrens–Fisher problem specifically, so I fully expect similarly good performance in many other important applications. Surely, the computational efficiency of my proposed solution can be improved, so I welcome advancements along these lines from computationally savvy readers.
The focus here has been on no-prior solutions, but there are cases in which incomplete or partial prior information is available, including high-dimensional problems in which structural assumptions like sparsity are common. Bayesian solutions can’t work with incomplete prior information—a prior distribution, perhaps diffuse, must be specified for every aspect of the unknown $\Theta$. Genuine partial-prior possibilistic IMs have recently been developed (e.g., Martin, 2022b ), and an interesting idea would be to apply the same inner probabilistic approximation proposed here in the no-prior case to the partial-prior case in the above references. Just as in this paper, these inner probabilistic approximations would enjoy some of the inherent reliability properties satisfied by the partial-prior IM, but the details remain to be investigated.
@@ Acknowledgments
The author thanks the Editor, Associate Editor, and two anonymous reviewers for their valuable comments on an earlier version of this paper. This work is partially supported by the U.S. National Science Foundation, grants SES–2051225 and DMS–2412628.
@@
$\sup_{\Theta\in\mathbb{T}}\mathsf{P}_{\Theta}\{\mathsf{\Pi}_{X}(H)\leq\alpha\text{ for some $H$ with $H\ni\Theta$}\}\leq\alpha,\quad\alpha\in[0,1].$ (17)
@@
There are mathematical obstacles that would make it difficult for any non-possibilistic quantification of uncertainty to satisfy the strong validity property in Equation (4) of the main paper, and hence the uniform calibration property in (17) above. Indeed, if $\mathsf{\Omega}_{x}$ was some other upper probability quantifying uncertainty about $\Theta$, given $X=x$, then one could define the analogue of its “contour,” i.e., $\theta\mapsto\mathsf{\Omega}_{x}(\{\theta\})$. But this may not meet the requirement of a possibility contour, that is, generally,
$\sup_{\theta\in\mathbb{T}}\mathsf{\Omega}_{x}(\{\theta\})<1,\quad\text{and often $\ll 1$}.$
Nothing in imprecise probability theory says that this “contour” must satisfy this supremum-equals-1 property and, in fact, in many (non-statistical) cases there are good reasons why such a property would be undesirable. But, as Shafer, (1976, p. 226) explains, statistical inference problems are special and well-suited for this supremum-equals-1 property. What I can add to Shafer’s justification is that it is virtually impossible for anything other than a possibilistic quantification of uncertainty, one for which the supremum-equals-1 property holds, to satisfy the validity property as stated in Equation (4) of the main paper and, therefore, the equivalent uniform calibration property in (17) above. The point is that, if $\mathsf{\Omega}_{X}(\{\Theta\})$ is bounded away from 1 for almost all $X$, then it can’t be stochastically no less than ${\sf Unif}(0,1)$. Moreover, if one was able to find a non-possibilistic IM $(\underline{\Omega}_{x},\mathsf{\Omega}_{x})$ that does satisfy the validity property in Equation (4) in the main text, then there is a possibilistic IM that is valid and no less efficient. So, there is at least no loss of generality in focusing on IMs that take on the possibilistic form.
@@
A point emphasized in Section 3.2 is that inner probabilistic approximations of a given IM $\mathsf{\Pi}_{x}$ aren’t unique. The reason is that the characterization result and the “maximally diffuse” objective don’t determine exactly how the kernel $\mathsf{K}_{x}^{\alpha}$ allocates its mass to the boundary $\partial C_{\alpha}(x)$. The goal of the present section is to expand a bit on this ambiguity to better understand the role played by the kernel.
To start, I’ll focus on the case of a scalar parameter, like in the illustrations presented in Section 3. In such a case, the kernel can be expressed generally as
where $w_{x,\alpha}$ is a data and level-dependent weight in $[0,1]$, $a(x,\alpha)$ and $b(x,\alpha)$ are the lower and upper endpoints of the confidence interval $C_{\alpha}(x)$, and $\delta_{z}$ is a probability distribution concentrating all of its mass on $z$. In this case, the kernel is fully determined up to the weight $w_{x,\alpha}$. My general recommendation was to take $w_{x,\alpha}$ to be constant equal to $\frac{1}{2}$, which has a two-fold motivation: it’s consistent with a version of the indifference principle and weights converging to $\frac{1}{2}$ is necessary in order for the inner probabilistic approximation to satisfy a Bernstein–von Mises theorem. Nevertheless, there’s still a choice of weight that can be made, so it’s worth having some understanding of what this entails.
For the gamma model $X\sim\mathsf{P}_{\Theta}={\sf Gamma}(n,\Theta)$ in Section 3.3 of the main text, Figure 2(a) shows the distribution functions associated with three different inner probabilistic approximations, each determined by a choice of the weight $w_{x,\alpha}$. Two of the three were based on constant-weight kernels: one with $w=0.5$ and one with $w\approx 0.45$; my rationale behind the choice $w\approx 0.45$ is described below. Since these weights are numerically very similar, it’s not surprising that the corresponding distribution functions are likewise similar. The third of the three distribution functions in Figure 2(a) is that based on the Bayes solution that employs the default right Haar prior with density $\theta\mapsto\theta^{-1}$ with respect to Lebesgue measure. I claimed in the main text that this Bayesian solution can’t be expressed in terms of the kernel above with constant weight; a non-constant weight $w_{x,\alpha}$ is needed. This begs the question: what non-constant weight corresponds to the right Haar prior Bayes solution?
In general, suppose that $Q_{x}(\theta)$ is the distribution function corresponding to an inner probabilistic approximation $\mathsf{Q}_{x}$; let $q_{x}$ be the corresponding density with respect to Lebesgue measure. Let $\hat{\theta}_{x}$ be the maximum likelihood estimator. Following the derivation in Section 3.3, for $\theta\leq\hat{\theta}_{x}$, the distribution function $Q_{x}(\theta)$ can be expressed as
$Q_{x}(\theta)=\int_{0}^{1}w_{\alpha}\,1\{a(x,\alpha)\geq\theta\}\,d\alpha=\int_{0}^{\pi_{x}(\theta)}w_{\alpha}\,d\alpha,$
where I’ve dropped the subscript “xx” in $w_{x,\alpha}$ for simplicity. Applying Leibniz’s integral rule gives the identity
$q_{x}(\theta)=w_{\pi_{x}(\theta)}\,\dot{\pi}_{x}(\theta),\quad\theta\leq\hat{\theta}_{x},$
where $\dot{\pi}_{x}(\theta)$ is the partial derivative of $\pi_{x}(\theta)$ with respect to $\theta$. This implies that
$w_{\alpha}=\frac{q_{x}(\theta_{x,\alpha})}{\dot{\pi}_{x}(\theta_{x,\alpha})},\quad\alpha\in[0,1],$ (18)
where $\theta_{x,\alpha}$ is such that $\theta_{x,\alpha}\leq\hat{\theta}_{x}$ and $\pi_{x}(\theta_{x,\alpha})=\alpha$. This calculation can’t be done analytically for the gamma illustration in Section 3.3, but it can be done numerically, and the plot of $\alpha\mapsto w_{\alpha}$ is shown in Figure 9. Notice that this curve ranges from below to above the $w=0.45$ line, with $w_{0.5}\approx 0.45$; this explains my choice of weight $w\approx 0.45$ as the best constant-weight approximation of the right Haar prior Bayes posterior.
Figure 9: Plot of $\alpha\mapsto w_{\alpha}$ as defined in (18) for the gamma model with $n=7$ and $x=14$.
The formula (18) for the non-constant weight is a bit abstract and can only be evaluated numerically in the gamma example above. So it may help to provide a simpler, proof-of-concept illustration—in particular, the Gaussian location model. In that case, the posterior density is $q_{x}(\theta)=(2\pi)^{-1/2}e^{-(\theta-x)^{2}/2}$, and the contour derivative is
$\dot{\pi}_{x}(\theta)=2(\theta-x)\times f_{1}((\theta-x)^{2}),$
where $f_{1}(z)=(2\pi)^{-1/2}z^{-1/2}e^{-z/2}$ is the ${\sf ChiSq}(1)$ density function. Then the ratio is
$\frac{q_{x}(\theta)}{\dot{\pi}_{x}(\theta)}=\frac{1}{2},\quad\theta<\hat{\theta}_{x}=x.$
This gives further justification for the choice of a constant weight $w=\frac{1}{2}$ in the IM’s inner probabilistic approximation’s kernel for the Gaussian location model.
More extreme choices of weights can be considered. For example, in the scalar parameter case, an extreme choice is taking $w=0$ or $w=1$. The large-sample results in Section 4.2 strongly suggest that such a choice isn’t warranted, so I wouldn’t recommend this. The point is that $w=0$ and $w=1$ both introduce bias that can’t be beneficial. When the parameter $\Theta$ is a vector or something more complex, then there are more choices of “extreme” kernels, some of which might not be entirely unreasonable. For instance, if $\Phi=m(\Theta)$ is a scalar feature of interest, then one might consider an “extreme” kernel that puts all of its weight on the extremes of the function $m$ on each confidence set boundary. That is, one might consider a kernel
$\mathsf{K}_{x}^{\alpha}=\tfrac{1}{2}\,\delta_{\underline{m}_{\alpha}}+\tfrac{1}{2}\,\delta_{\overline{m}_{\alpha}},$
where
$\underline{m}_{\alpha}=\min_{\theta\in\partial C_{\alpha}(x)}m(\theta)\quad\text{and}\quad\overline{m}_{\alpha}=\max_{\theta\in\partial C_{\alpha}(x)}m(\theta).$
With this choice, the marginal distribution $\mathsf{Q}_{x}^{m}$ of $\Phi=m(\Theta)$ derived from the inner probabilistic approximation with the above kernel is my recommended inner probabilistic approximation of the marginal IM $\mathsf{\Pi}_{x}^{m}$ derived from $\Pi_{x}$ via extension. The point is that “extreme choices” of kernels might not be unreasonable, and the justification of such a choice would be based on marginalization considerations, when there’s a particular feature $\Phi=m(\Theta)$ of $\Theta$ of interest. When there’s no single feature that’s of high priority, it makes sense to make a neutral choice of kernel, and, aside from the large-sample properties, it’s precisely the uniform kernel’s neutrality that makes it a good choice.
Appendix C Details from Section 4.1
C.1 Theory
@@
Recall that $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ is a family of probability distributions supported on $\mathbb{X}$ that is also invariant with respect to a group of transformations $\mathscr{G}$ from $\mathbb{X}$ to itself in the sense of Equation (10) in the main text. If the distributions $\mathsf{P}_{\theta}$ have densities with respect to some underlying dominating measure, which I’ll assume here, then invariance means
@@
where $\chi(g)$ is the “multiplier,” a change-of-variables Jacobian term; see Schervish, (1995, Ch. 6) and Eaton, (1989) for details. A further simplification that I’ll adopt here, common in the literature (e.g., Schervish, 1995, p. 371), is to assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$. Recall that I’m also assuming here that the group $\mathscr{G}$ with respect to which the model is invariant is unique. There are cases, however, where the model is invariant with respect to more than one group, and I’ll elaborate on this case at the end of this subsection.
@@
Suppose that the model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ is invariant with respect to a (unique) group $\mathscr{G}$, and that the corresponding density functions $\{p_{\theta}:\theta\in\mathbb{T}\}$ satisfy (19). For simplicity, assume $\mathbb{T}=\mathscr{G}=\overline{\mathscr{G}}$. In addition, assume that the following hold:
@@
A couple quick remarks are in order. First, for A1, existence and uniqueness of the left and associated right Haar measures is a classical result, which holds under the mild condition that $\mathscr{G}$ be a locally compact topological group; see Halmos, (1950) and Nachbin, (1965). For A3, existence of a joint density with respect to a product measure ensures that there’s no difficulty in defining a conditional distribution for $G$, given $U=u$, which is needed below. Finally, $U=U_{X}$ is an ancillary statistic as a function of $X\sim\mathsf{P}_{\theta}$, meaning that its marginal distribution doesn’t depend on $\theta$.
The simplest example is a location model where $\mathscr{G}=\overline{\mathscr{G}}=(\mathbb{R},+)$. This is an abelian group, so the left and right Haar measures are the same and both equal to Lebesgue measure. The function $x\mapsto t(x)$ in A2 consists of two components: in its $g_{x}$” coordinate an equivariant function of $x$ that estimates the location and, in its $u_{x}$” component, an invariant function of $x$, such as residuals. For example, $g_{x}=\bar{x}$ the arithmetic mean of $x=(x_{1},\ldots,x_{n})$ and $u_{x}=\{x_{i}-\bar{x}:i=1,\ldots,n\}$. Note that the $u_{x}$ coordinate satisfies a constraint, so, after it’s represented in a suitable lower-dimensional space $\mathbb{U}$, $\mu$ can be taken as Lebesgue measure there. There are many other problems that fit this general form; see, e.g., Chapters 1–2 of Fraser, (1968), including the exercises.
@@
  • Lemma 6.65 in Schervish, (1995) gives a Bayesian posterior distribution $\mathsf{Q}_{x}^{\rho}$ for $\Theta$ that has the following density with respect to the right Haar prior measure $\rho$:
    $q_{x}^{\rho}(\theta)=\frac{d\mathsf{Q}_{x}^{\rho}}{d\rho}(\theta)\propto f(\theta^{-1}\circ g,u),\quad\theta\in\mathbb{T},\quad x\equiv(g,u).$ (21)
@@
The above points are crucial to the following calculation. Indeed, since the $U$ component of $X$ is ancillary, the minimum complexity principle in Martin, 2022b says the IM construction should condition on the observation $U=u$. Write $\pi_{g|u}(\theta)$ for the possibilistic IM’s contour, where the subscript is meant to emphasize that the conditional distribution of $X\equiv(G,U)$, given the observed value of $u$, is used in the validification step (2). Then the IM contour is defined as
$\pi_{g|u}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\mid U=u\},\quad\theta\in\mathbb{T}.$ (22)
Then the following identities are consequences of the two results above that link the relative likelihood and the posterior density:
@@
$=\mathsf{Q}_{x}^{\rho}\{f(H,u)\leq f(\theta^{-1}\circ g,u)\}$
$=\mathsf{Q}_{x}^{\rho}\{q_{x}^{\rho}(\Theta)\leq q_{x}^{\rho}(\theta)\},$
@@
where the first line is by simplification of the definition of $\pi_{g|u}$; the second line is by the fact that $H:=\theta^{-1}\circ G$ is a pivot with respect to the conditional distribution of $G$, given $U=u$, under $\mathsf{P}_{\theta}$ (e.g., Schervish, 1995, Cor. 6.67); the third line is by the fact that $H:=\Theta^{-1}\circ g$ is also a pivot with respect to the Bayes posterior $\Theta\sim\mathsf{Q}_{x}^{\rho}$ and has the same distribution as $H$ in previous line; and the last line is by simplification via (21).